# Quantum field theory on curved manifolds

### v1

@@ 3 The Schwinger effect on manifolds
First, we consider the case where the curvature of the manifold is defined for a gauge symmetry. The simplest model uses the electromagnetic $U(1)$ gauge symmetry on a flat space-time and the curvature is introduced by a constant electric field. In this model, the manifold is static in the sense that the curvature is constant, but quantum theory expects a dynamical phenomenon on it. The particle production in this model is called the Schwinger effect[1]. In this case, the quickest way to avoid tedious discussions about manifolds is to use a powerful computational tool, the path integral[1, 10, 11]. However, in this paper, we venture a primitive analysis based on field equations to look closely at what happens on the manifold. The analysis on the manifold using a scalar field is already given in Ref.[12, 13, 14]. What is new in this paper is the analysis of the fermionic Schwinger effect as the Landau-Zener transition and its application to a time-dependent electric field. We show that the concept of the manifold is particularly important when considering a slowly varying electric field.
@@
Cosmological particle production after inflation has been discussed for time-dependent mass $m(t)$ in various situations. Such particle production is called preheating[18, 19]. It was first recognized in Ref.[20] that the Fermion preheating can be interpreted as the Landau-Zener transition[8], and the idea has been extended in Ref.[16, 17] to solve cosmological problems of particle production. Particle-antiparticle asymmetry is not discussed here, but when it is described by the multi-element Landau-Zener transition, there are seeds of asymmetry in the interference between different kinds of the Stokes phenomena[16, 17]. The relationship between cosmological particle production and the Landau-Zener transition is not discussed in detail in this paper, so more details and further explanations are left to these papers.
@@ 3.1 Constant electric field (constant curvature)
Decoupling the equations, we have [^8] One might claim that the equation can be solved immediately using special functions. The reason for the somewhat roundabout approach here is that we want to examine the Stokes lines in the vicinity of the tangent space. To understand the structure of the Stokes lines, we use the Exact WKB (EWKB) developed in Refs.[21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32].
@@
$\ddot{\hat{X}}$ $+$ $\frac{|\Delta|^{2}+D^{2}}{\hbar^{2}}\hat{X}=0.$ (3.33)
When $\Delta(t)=\lambda t$ and $D(t)=D_{0}$, the above equation is a well-known problem of quantum mechanics (i.e, scattering by an inverted quadratic potential).
@@
There are a few things to be considered with care when using this equation for solving the Schwinger effect. The first and the most important is the definition of the vacuum. In the usual analysis of a constant electric field, one will find a scattering problem by an quadratic potential, where the two vacuum states (in and out states at $t=-\infty$ and $t=+\infty$ where the electric field is supposed to disappear) are defined as asymptotic states. Then, the Stokes phenomenon is assumed to occur between the two vacuum states. In this case, the gauge symmetry is used to explain the arbitrariness of positioning the top of the potential hill. However, as already explained, tangent spaces are naturally introduced in manifolds, and these tangent spaces have the property for defining the local vacuum states. Therefore, instead of losing the locality of the analysis by assuming the vacuum states at far away, we try to solve the problem by defining the vacuum on the tangent space. By definition, a feature of the tangent space is that the connection is zero in the space. This feature must be realized on the manifold at the point of contact. This is also a simple consequence of the local trivialization. This means that the gauge symmetry of each $U_{i}$ must be used carefully to make the connection vanish at the point. This makes the top of the quadratic potential coincide with the tangent space, as is illustrated in Fig.1.
@@
Figure 1: The typical “potential” of the equation is shown for $V(t)=1-t^{2}$. The Stokes lines for the potential are shown on the complex $t$-plane. They are presented on the top. The Stokes lines appearing on $U_{i}$ of $M$ are illustrated in the bottom picture.
@@
Then, they are laminated using gauge transformations. The situation is quite similar to the moving frame in special relativity. Similarly, since the tangent space is also a local inertial system, the velocity (or $k$) is assumed to be negligible to account for the inertial vacuum seen by the generated particle. [^10] In the Unruh effect, the vacuum is non-trivial when it is seen by the particle, not by someone else. Note that choosing a section of the observer (i.e, the inertial frame) does not kill the Lorentz symmetry of the theory. Similarly, the choice of the gauge in $U_{i}$ does not kill the gauge symmetry of the theory. This changes the definition of the electric field $E_{0}$ in the equation since the original $E_{0}$ was defined for an observer in the laboratory. Therefore, we denote the electric field in the equation described in the vicinity of the tangent space as $\hat{E}$ in the following calculations. Now we have defined everything in the vicinity of the contact point of the tangent space. Note that the “decoupling” of the equations performed above cannot be defined at the point of contact with the tangent space, as the non-diagonal element disappears from the matrix and the equations are already decoupled there. This shows that there is no mixing in the tangent space by definition. As discussed in the definition of manifolds, the Stokes phenomenon should be considered on an open set $U_{i}$ defined in a neighborhood of $p\in M$, and it is the physics around $p$ that is relevant for mixing solutions. Therefore, the Stokes phenomenon is considered here in a neighborhood where the point $\Delta=0$ is slightly avoided. The above equations show that the real-time axis traverses the Stokes line in the vicinity of the tangent space on every $U_{i}$[13]. See also Fig.1.
@@
Thus, when the local setting is made natural as a manifold, it can be seen that the Stokes lines appear in the neighborhood of the local tangent space without ambiguity of the time and the gauge. This indicates that the Stokes Phenomenon is constantly mixing the vacuum solutions. Exactly speaking, the vacuum defined “in” the tangent space cannot describe the mixing by definition, while the mixing is seen in the neighborhood of the tangent space placed on each $U_{i}$. This is the same situation as the asymptotic states. In the same way as the correspondence with the vacuum is considered for the asymptotic states, the correspondence with the vacuum solution here is considered in the vicinity of the tangent space.
@@
Here, one might notice that the frame prepared for the generated particles of the Schwinger effect actually represents an accelerated frame called the moving frame. For this simple reason, the analysis of the Schwinger effect would not be complete without an analysis of the Unruh effect. This topic will be discussed in the next section. In the remaining part of this section, we are going to look at the time-dependent case in a little more detail.
@@ 3.2 Time-dependent $E(t)$
We will now try to show an example where it makes no sense at all as physics to find a naive global solution to an equation defined on an open set $U_{i}$. It naturally depend on the situations whether a meaningful result can be obtained by such solutions when extrapolating the equations originally defined on $U_{i}$. However, almost all papers do not mention this point at all. In order to understand the problem without ambiguity, consider the case where the electric field changes gradually and there is no significant back-reaction from particle production. If the electric field oscillates, our assumption here is that the time period of the oscillation is much longer than the width of $U_{i}$. [^11] If the period of the oscillation is short, averaging over a period should be considered for the observation. In addition, the Stokes phenomenon in mathematics occurs on the Stokes “line” on the complex plane, but in physics, it will be better to assume an intrinsic width of the line. Therefore, the width of on an open set $U_{i}$ is assumed to be wider than the intrinsic width of the Stokes line.
Replacing $E_{z}(t)=E_{0}$ with $E_{z}(t)=\alpha t$, the problem becomes scattering by a “quartic” potential. The solution to this problem and the Stokes lines have been studied in great detail by Voros[22], and it has been found that the Stokes lines act away from the origin. See Fig.2 for more details about the Stokes lines. Details concerning conventional particle production can be found in Ref.[20]. Unlike the case of the constant electric field, the gauge ambiguity cannot naively shift the position of the quartic potential. Also, although intuitively the probability of particle production should change gradually if the electric field changes slowly, such calculation (scattering by the quartic potential solved for asymptotic states) will not change in such a way. Although it is immediately obvious to anyone who looks at the equation, we point out that the solutions of the equation, if they are used for the asymptotic states, can not solve the problem at all. It is easy to say that the setting of the problem is bad, but here we are going to try to clarify our understanding a little bit more. For our purpuses, it is important to go back to the definition of the manifold and solve the problem using local analysis.
@@
Let us now reconsider the above question along the lines of the manifold and the differential geometry. First, consider an open set $U_{i}$ in the neighborhood of $t=t_{p}$ and let $A_{z}(t)=\frac{1}{2}\alpha t^{2}$ be expanded at $t=t_{p}$. Then, we have the “local potential” defined for $U_{i}$ as
$V(t)|_{U_{i}}$ $\simeq$ $e^{2}\hat{E}_{z}(t_{p})^{2}(t-t_{p})^{2},$ (3.34)
where by using the gauge of $U_{i}$ the connection is set to zero at the contact point $t=t_{p}$. This equation clearly meets the above requirement. In this way, the generation rate can be calculated for the local time ($t_{p}$) on the local open set ($U_{i}$).
There could be critical opinions about not defining asymptotic states for solving field equations to analyze particle production, but at least for cases like the one discussed here, the use of the tangent space seems to be inevitable. We would like to stress the importance of using two different definitions of the vacuum for different purposes. Furthermore, if the electric field changes much more rapidly and is treated as quantized, then the above treatment as a “background field” would no longer be appropriate and the field equations of the gauge field come into play. Then, we cannot imagine a better alternative to Feynman diagrams and path integrals.
@@
As we have already seen in the simple example of a slowly changing electric field, it is very important and sometimes quite essential to solve the field equations locally on the manifold. Of course, when determining the averaged particle production rate in the case of repeated rapid oscillations or when the electric field is instantaneous, local evaluation of the particle production rate is useless for the experimental observation. Then, to get the required results, one should consider asymptotic states and averaging over a period. [^12] Analysis of these topics using the Stokes lines can be found in Ref.[33]. See also Refs.[34, 35].. What we have highlighted in this section is the case in which local calculation is needed to obtain correct results while a “naive” application of the conventional asymptotic states gives clearly a wrong answer. This distinction has never been explicitly recognized. We believe that the importance of the option of defining the vacuum in the tangent space has been made very clear by the simple model described in this section.
Let us now look at how such a definition of the vacuum has implications in relativity.
@@ 4 The Unruh effect and Hawking radiation on manifolds
The previous section dealt with the case where there is no qualitative ambiguity in defining the local vacuum, but in general relativity, even after the connection and the tangent spaces are determined, there are still further degrees of freedom left in the vierbein, which leaves ambiguity in defining the local vacuum. The most obvious difference is between the Lorenz frame and the local inertial frame. In mathematics, covariant derivatives are defined by using the Lorenz frame, which gives a vierbein that is diagonal with respect to the time-direction in the neighborhood. On the other hand, in physics, the vacuum is defined for the local inertial frame, which gives a vierbein that is diagonal with respect to the time-direction only at the contact point. In both cases, the tangent space is correctly defined at the point, but there is a difference in physics in the neighborhood.
@@
What is important for our discussion is that in an $m$-dimensional Riemannian manifold, the metric tensor $\mathrm{g}_{\mu\nu}$ has $m(m+1)/2$ degrees of freedom while the vielbein has $m^{2}$ degrees of freedom. For $m=4$, we have $10$ for the metric while $16$ for the vielbein. They are not identical. Each of the bases can be related to the other by the local orthogonal rotation $SO(m)$, while for the Lorentzian manifold it becomes $SO(m-1,1)$. The dimension of these Lie groups is given by the difference between the degrees of freedom of the vielbein and the metric. This shortly means that there are many (uncountable) choices for non-coordinate bases even after the metric is identified. This point will be very important when one looks at the Unruh effect[36, 37, 7]. The local inertial frame and the Lorentz frame have the same metric and are defining the same tangent space at the point. However, they are distinguished by the vierbein. The different choices of the “vacuum” is essential in the search for the Stokes phenomenon[13].
We describe the frame bundle further below. Associated with a tangent bundle $TM$ over $M$ is a principal bundle called the frame bundle $LM\equiv\bigcup_{p\in M}L_{p}M$ where $L_{p}M$ is the set of frames at $p$. The bundle $T_{p}M$ has a natural coordinate basis $\{\partial/\partial x^{\mu}\}$ on $U_{i}$ and a “frame” $u=\{X_{1},...,X_{m}\}$ at $p$ is expressed by the non-coordinate basis
@@
The following point is very important for our discussion. A natural coordinate basis is prepared on the surface of $M$ and the inertial system is defined using a non-coordinate basis. This procedure naturally gives a notion of the “moving frame”[9], as the inertial frame seems to be rotated from time to time by the Lorentz transitions of the vielbeins, somewhat like spinning tea cups in amusement parks. Thomas precession is explained by the fact that multiple Lorentz transformations with different directions produce a rotation of the intrinsic space of the observer. If the observer stays in the same frame (no acceleration), there is no motion in the direction of the fibre of the frame bundle. In this case, the connection vanishes by definition because there is no transformation when laminating. Therefore, for an inertial observer, the distinction between coordinate and non-coordinate systems will be quite ambiguous. However, if one wants to describe an observer in accelerated motion on flat space-time, one has to define the local inertial frame for the observer, which moves on the frame bundle in a non-trivial way. This defines the section of the observer on the frame bundle. This means that on this section, one has to laminate the bundle by using non-trivial coordinate transformations. This (the observer’s non-trivial section of the frame bundle) introduces the connection for the observer, although the space-time is flat.
@@
Indeed, for such constant acceleration, the vielbeins have completely the same form for any time $(\tau)$. The transformation $(e_{A})_{\alpha}^{\mu}\rightarrow(e_{B})_{\alpha}^{\mu}$ on the frame bundle is the Lorentz transformation;
@@
Note that the situation is very similar to the Schwinger effect for a constant electric field. In the Schwinger effect, the potential has been shifted by a gauge transformation; in the Unruh effect, the Lorentz transformation gives vielbeins of exactly the same shape. In both cases, the observer is always looking at the same physics on a static manifold.
Of course, the mathematical description of the frame bundle does not require an observer, but in physics an observer is inevitable. In this case, the observer is nothing but a section of the frame bundle. Previously, we have introduced $d\phi$ to describe covariant derivatives of the gauge symmetry. We have seen that both $\phi$ and $d\phi$ are the sections of the vector bundle, and the connection is introduced to make $d\phi$ covariant by the gauge symmetry. Now we can see that $d\phi$ is also a section of the frame bundle. If an observer stays in the same frame, there is no (non-trivial) connection required for $d\phi$. However, when an observer is accelerating, $d\phi$ traverses the frame bundle in the direction of the fibre and its motion causes Lorentz transformation. This is called “the moving frame”.
@@
Is local analysis possible also for the Unruh effect, if we follow the previous calculations of the Schwinger effect? Our answer is “No”. What is important here is that the definition of the covariant derivative uses the Lorenz frame in which the vierbein is diagonalized in the neighborhood. On the other hand, as is shown explicitly above, the vierbein of the local inertial frame is diagonalized only at the point. (Note that the off-diagonal elements vanish since $\sinh a(\tau-\tau_{A})=0$ at $\tau=\tau_{A}$.) Since the covariant derivatives are defined for the Lorentz frame while the Unruh effect is defined for the inertial frame, it seems impossible to examine the Stokes phenomena of the Unruh effect directly (and locally) in terms of the field equations[13]. This mismatch has prevented the local analysis of the Unruh effect for a long time.
@@
Now consider the physics that observers see in the Unruh effect. As far as the acceleration seen by the observer is constant, the physics seen by the observer is indistinguishable at any time due to the equivalence classes defined for the manifold. (More particularly, the observer feels the same vierbein all the time.) This is purely a mathematical consequence. If a dynamical effect (particle production) is manifested in such a situation, it must be explained by the vierbein of the local inertial frame. Extrapolating the coordinates to infinity and considering a global map as the Bogoliubov transformation is conceptually unacceptable, even if it could be plausible as a method[3]. In fact, it is known that such methods can lead to unnatural entanglements appearing between regions that should be uncorrelated. To argue the legitimacy of our local computation on manifolds, we must address this issue in this paper.
@@
First of all, consider what happens if the scalar field equations were written down for an accelerating observer. Rindler coordinates are a specific set of coordinates used in the context of special relativity to describe the motion of an observer undergoing constant proper acceleration in flat spacetime. Thus, the Rindler coordinates form a coordinate chart that covers a specific region of Minkowski spacetime known as the Rindler wedge. Following the concept of manifolds, the simplest local inertial system is defined as a tangent space in the neighborhood of the point with zero velocity in Rindler’s coordinates. We start with the conventional Rindler metric[3, 9, 38];

### v5

@@ 3 The Schwinger effect on manifolds
First, we consider the case where the curvature of the manifold is defined for a gauge symmetry. The simplest model uses the electromagnetic $U(1)$ gauge symmetry on a flat space-time and the curvature is introduced by a constant electric field. In this model, the manifold is static in the sense that the curvature is constant, but quantum theory expects a dynamical phenomenon (particle production) on it. The particle production in this model is called the Schwinger effect[3]. In this case, the quickest way to avoid tedious discussions about manifolds is to use a powerful computational tool, the path integral[3, 13, 14]. However, in this paper, we venture a primitive analysis based on field equations to look closely at what happens in differential geometry and manifolds. The analysis on the manifold using a scalar field is already given in Ref.[15, 16, 17]. The analysis of the fermionic Schwinger effect as the Landau-Zener transition and its application to a time-dependent electric field is new. Using this model and a slowly varying electric field, we show why the concept of the manifold and definition of local vacuum in the tangent space is particularly important.
@@
Cosmological particle production after inflation has been discussed for time-dependent mass $m(t)$ in various situations. Such particle production is called preheating[21, 22], for which there is no serious need for careful discussions about locality. A serious discussion about locality is important when the system appears to be static and a symmetry is related to it. The most important topic in this section is how this relationship between locality and symmetry is described in differential geometry.
It was first recognized in Ref.[23] that the Fermion preheating can be interpreted as the Landau-Zener transition[11], and the idea has been extended in Ref.[19, 20] to solve cosmological problems of particle production. Particle-antiparticle asymmetry is not discussed here, but when it is described by the multi-element Landau-Zener transition, there are seeds of asymmetry in the interference between different kinds of the Stokes phenomena[19, 20]. The relationship between cosmological particle production and the Landau-Zener transition is not discussed in detail in this paper, so more details and further explanations are left to these papers.
@@ 3.1 Constant electric field (constant curvature)
Decoupling the equations, we have [^11] One might claim that the equation can be solved immediately using special functions. The reason for the somewhat roundabout approach here is that we want to examine the Stokes lines in the vicinity of the tangent space. To understand the structure of the Stokes lines, we use the exact WKB developed in Refs.[24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35].
@@
$\ddot{\hat{X}}$ $+$ $\frac{|\Delta|^{2}+D^{2}}{\hbar^{2}}\hat{X}=0,$ (3.33)
where $V(t)\equiv-(|\Delta|^{2}+D^{2})$ is called a “potential” of the “Schrodinger equation”. When $\Delta(t)=\lambda t$ and $D(t)=D_{0}$, the above equation is similar to a well-known problem of quantum mechanics (i.e, the scattering by an inverted quadratic potential). The solutions of $|\Delta(t)|^{2}+D(t)^{2}=0$ are called “turning points”. The main difference between the analysis using the exact WKB and the conventional analysis is that the local analysis in the vicinity of the Stokes line is mathematically well defined in the exact WKB. This distinguishing property of the exact WKB is the most important when considering local analysis.
@@
There are a few things to be considered with care when using this equation for solving the Schwinger effect. The first and the most important is the definition of the local vacuum. In the usual analysis of a constant electric field, one will find a scattering problem by an quadratic potential, where the two vacuum states (in and out states at $t=-\infty$ and $t=+\infty$ where the electric field is supposed to disappear) are defined as asymptotic states. Then, the Stokes phenomenon is assumed to occur between the two vacuum states. In this case, the gauge symmetry is used to explain the arbitrariness of positioning the top of the potential hill. However, as already explained, tangent spaces are naturally introduced in manifolds, and these tangent spaces have the property for defining the local vacuum states. Therefore, instead of losing the locality of the analysis by assuming the vacuum states at far away, we try to solve the problem by defining the vacuum in the tangent space. As we have already mentioned, a feature of the tangent space is that the curvatur is zero in the space, and the connection must vanish at the point of contact. This is also a simple consequence of the local trivialization we have described above. This means that the gauge symmetry of each $U_{i}\in M$ must be used carefully to make the connection vanish at the point. Although it may not be trivial, one can see from the equation that this makes the top of the quadratic potential coincide with the tangent space in each $U_{i}$, as is illustrated in Fig.3.
@@
Figure 3: The typical “potential” of the equation is shown for $V(t)=1-t^{2}$. The Stokes lines for the potential are shown on the complex $t$-plane. They are presented on the top. The Stokes lines appearing on $U_{i}$ of $M$ are illustrated in the bottom picture. The top of the local “potential” must coincide with the contact point of the tangent space because of the local trivialization in each $U_{i}$. This makes the Stokes lines coincide at the contact point and makes each $U_{i}$ looks completely the same.
@@
Then, they are laminated using gauge transformations. The situation is quite similar to the moving frame in special relativity. Similarly, since the tangent space is also a local inertial system, the velocity (or $k$) is assumed to be negligible to account for the inertial vacuum seen by the generated particle. This changes the definition of the electric field $E_{0}$ in the equation since the original $E_{0}$ was defined for an observer in the laboratory. Therefore, we denote the electric field in the equation described in the vicinity of the tangent space as $\hat{E}$ in the following calculations. Now we have defined everything in the vicinity of the contact point of the tangent space. Note that the “decoupling” of the equations performed above cannot be defined at the point of contact with the tangent space, as the non-diagonal elements disappear from the matrix and the equations are already decoupled there. This shows that there is no mixing in the tangent space by definition. As discussed in the definition of manifolds, the Stokes phenomenon should be considered on an open set $U_{i}$ defined in a neighborhood of $p\in M$, and it is the physics around $p$ that is relevant for mixing solutions. Therefore, the Stokes phenomenon is considered here in a neighborhood. The above equations show that the real-time axis traverses the Stokes line in the vicinity of the tangent space on every $U_{i}$[16]. See also Fig.3.
@@
Thus, when the local setting is made natural as a manifold, it can be seen that the Stokes lines appear in the neighborhood of the local tangent space without ambiguity of the time and the gauge. This indicates that the Stokes Phenomenon is constantly mixing the vacuum solutions. Exactly speaking, the vacuum defined “in” the tangent space cannot describe the mixing by definition, while the mixing is seen in the neighborhood of the tangent space placed on each $U_{i}$. This is the same situation as the asymptotic states. In the same way as the correspondence with the vacuum is considered for the asymptotic states, the correspondence with the vacuum solutions here is considered in the vicinity of the tangent space.
@@
Here, one might notice that the frame prepared for the generated particles of the Schwinger effect actually represents an accelerated frame called the moving frame. For this simple reason, the analysis of the Schwinger effect would not be complete without an analysis of the Unruh effect. This topic will be discussed in the next section. In the remaining part of this section, we are going to look at the time-dependent case in a little more detail to show a simple example in which definition of the local vacuum is crucial for the calculation.
@@ 3.2 Time-dependent $E(t)$
We will now try to show an example where it makes no sense at all as physics to find a naive global solution by considering extrapolation to an equation defined on an open set $U_{i}$. It naturally depend on the situations whether a meaningful result can be obtained by such extrapolation. However, as far as we know, there has been no paper in which this point is mentioned. In order to understand the problem without ambiguity, consider the case where the electric field changes gradually and there is no significant back-reaction from particle production. If the electric field oscillates, our assumption here is that the time period of the oscillation is much longer than the width of $U_{i}$.
Replacing the constant electric field $E_{z}(t)=E_{0}$ with a slowly varying electric field $E_{z}(t)=\alpha t$, the problem becomes scattering by a “quartic” potential $V(t)\propto-t^{4}$. The solution to this problem and the Stokes lines have been studied in great detail by Voros[25], and it has been found that the Stokes lines act away from the origin. See Fig.4 for more details about the Stokes lines. Details concerning conventional particle production can be found in Ref.[23]. What is important here is that unlike the case of the constant electric field, the gauge ambiguity cannot naively shift the position of the quartic potential. Also, although intuitively the probability of particle production should change gradually if the electric field changes slowly, a conventional calculation (scattering by the quartic potential solved for asymptotic states) will not change in such a way. Although it shoule be immediately obvious to anyone who looks at the equation, we point out that the solutions of the equation, if they are used for the asymptotic states, can not solve the problem at all, even if the solutions are exact. It is easy to finish the discussion by saying that the setting of the problem is bad, but here we are going to try to clarify our understanding a little bit more. For our purpuses, it is important to go back to the definition of the manifold and solve the problem in local.
@@
Let us now rephrase the above question along the lines of the manifold and the differential geometry. First, consider an open set $U_{i}$ in the neighborhood of $t=t_{p}$ and let $A_{z}(t)=\frac{1}{2}\alpha t^{2}$ be expanded at $t=t_{p}$. Then, we have the “local potential” defined for $U_{i}$ as
$V(t)|_{U_{i}}$ $\simeq$ $m^{2}-e^{2}\hat{E}_{z}(t_{p})^{2}(t-t_{p})^{2},$ (3.34)
where by using the gauge of $U_{i}$ the connection ($A_{\mu}$) is set to zero at the contact point $t=t_{p}$. $k\simeq 0$ is also expected for the inertial frame of the particle. Here the tangent space and the local vacuum is defined at $t=t_{p}$. This equation clearly meets the above requirement if the local vacuum is defined in the tangent space. See also Fig.5. The particle production can be calculated locally on each $U_{i}\in M$ by using $V(t)|_{U_{i}}$ and the production rate gradually changes with time. In this way, the generation rate can be calculated for the local time ($t_{p}$) on the local open set ($U_{i}$). For cases like the one discussed here, the use of the tangent space to define the local vacuum is inevitable. The conventional calculation defining the asymptotic vacuum at a distance could be convenient but not rigorous in mathematics. We would like to stress the importance of understanding benetits and risks of the definitions of the vacuum. As far as we can understand, the serious risk of defining the vacuum in the asymptotic state has not been explained in detail so far for the Schwinger effect.
Figure 5: The upper picture shows the Stokes lines of the solution when the equation defined at $t=0$ is extrapolated to infinity. The lower picture shows the Stokes lines when equations are defined for each $U_{i}$ using local trivialization. One can see that the production rate is gradually changing with time (as the distance between two turning points are changing with time).
@@
As we have already seen in the simple example of a slowly changing electric field, it is very important and sometimes quite essential to solve the field equations locally on the manifold by defining a local vacuum in the tangent space. Of course, when determining the averaged particle production rate in the case of rapid oscillations or when the electric field is sharply instantaneous, local evaluation of the particle production rate should be useless for experimental observation. If the electric field changes rapidly, the field equation becomes a simultaneous differential equation with the electromagnetic field equation. The Stokes phenomena in such cases can be extremely complex and require appropriate approximations. [^13] Analysis of these topics using the Stokes lines can be found in Ref.[36]. See also Refs.[37, 38].. What we have highlighted in this section is the case in which local calculation is crucial to obtain correct results while a “naive” extrapolation of the equation to the outside of the defined area gives clearly a wrong answer. This distinction has never been explicitly recognized so far. We believe that the importance of the option of defining the local vacuum in the tangent space has been made very clear by the simple model described in this section.
Let us now look at how such a definition of the vacuum has implications in relativity. As mentioned above, particles created by the Schwinger effect can also be affected by the Unruh effect. In this case, the question is whether the two effects are independent phenomena or whether we are simply seeing the same thing. In the following, we present the solution to this problem using differential geometry and the exact WKB.
@@ 4 The Unruh effect and Hawking radiation on manifolds
The previous section dealt with the case where there is no qualitative ambiguity in defining the local vacuum, but in general relativity, even after the connection (metric) is determined, there are still further degrees of freedom left in the vierbein, which leaves ambiguity in defining the local vacuum. The most obvious difference is between the Lorenz frame and the local inertial frame. In mathematics, covariant derivatives are defined by using the Lorenz frame, which gives a vierbein that is diagonal with respect to the time-direction in the neighborhood. On the other hand, in physics, the vacuum is defined for the local inertial frame, which gives a vierbein that is diagonal with respect to the time-direction only at the contact point. In both cases, the tangent space is correctly defined at the point, but there is a difference in physics in the neighborhood.
@@
What is important for our discussion is that in an $m$-dimensional Riemannian manifold, the metric tensor $\mathrm{g}_{\mu\nu}$ has $m(m+1)/2$ degrees of freedom while the vielbein has $m^{2}$ degrees of freedom. For $m=4$, we have $10$ for the metric while $16$ for the vielbein. They are not identical. Each of the bases can be related to the other by the local orthogonal rotation $SO(m)$, while for the Lorentzian manifold it becomes $SO(m-1,1)$. The dimension of these Lie groups is given by the difference between the degrees of freedom of the vielbein and the metric. This shortly means that there are many (uncountable) choices for non-coordinate bases even after the metric is identified. This point will be very important when one looks at the Unruh effect[39, 40, 5]. The local inertial frame and the Lorentz frame have the same metric and are defining the same tangent space at the point. However, they are distinguished by the vierbein. The difference in the vierbein is essential in the search for the Stokes phenomenon of the Unruh effect[16].
We describe the frame bundle further below. Associated with a tangent bundle $TM$ over $M$ is a principal bundle called the frame bundle $LM\equiv\bigcup_{p\in M}L_{p}M$ where $L_{p}M$ is the set of frames at $p$. Here we have a natural coordinate basis $\{\partial/\partial x^{\mu}\}$ and a “frame” $u=\{X_{1},...,X_{m}\}$ at $p$ is expressed by the non-coordinate basis
@@
The following point is very important for our discussion. A natural coordinate basis is prepared for $U_{i}\in M$ and the inertial system is defined using a non-coordinate basis in the tangent space. This procedure naturally gives a notion of the “moving frame”[12], as the inertial frame seems to be rotated from time to time by the Lorentz transitions of the vielbeins, somewhat like spinning tea cups in amusement parks. [^15] One can see more explanations in Ref.[12], in which figures for the moving frame can be found. The Thomas precession is explained by the fact that multiple Lorentz transformations with different directions produce a rotation of the intrinsic space of the observer. If the observer stays in the same frame (no acceleration), there is no motion in the direction of the fibre of the frame bundle. In this case, the connection vanishes by definition. Therefore, for an inertial observer, the distinction between coordinate and non-coordinate systems will be quite ambiguous. However, if one wants to describe an observer in accelerated motion on flat space-time, one has to define the local and subjective inertial frame for the observer in the tangent space, which moves on the frame bundle in a non-trivial way. This defines the section of the observer on the frame bundle. This means that on this section, one has to laminate the bundle by using non-trivial coordinate transformations. The Thomas precession is explained that the lamination by the Lorentz transformation causes rotation of the intrinsic space. This (the observer’s non-trivial section of the frame bundle) introduces the connection for the observer, although the space-time is flat.
@@
Indeed, for such constant acceleration, the vielbeins have completely the same form for any time $(\tau)$. At $\tau=\tau_{B}$, we have
$(e_{B})_{\alpha}^{\mu}$ $=$ $\left(\begin{array}[]{cc}\cosh a(\tau-\tau_{B})&\sinh a(\tau-\tau_{B})\\ \sinh a(\tau-\tau_{B})&\cosh a(\tau-\tau_{B})\end{array}\right).$ (4.11)
The transformation $(e_{A})_{\alpha}^{\mu}\rightarrow(e_{B})_{\alpha}^{\mu}$ on the frame bundle is the Lorentz transformation;
@@
Note that the situation is very similar to the Schwinger effect for a constant electric field. In the Schwinger effect, the potential has been shifted by gauge transformation and the equation looks always the same on each $U_{i}$. In the Unruh effect, the Lorentz transformation gives vielbeins of exactly the same shape on each $U_{i}$. In both cases, the observer is always looking at the same physics on a static manifold.
@@
Our primary question here is “Is the same local analysis possible also for the Unruh effect, if we follow the previous calculations of the Schwinger effect?” Our answer is “No”. What is important here is that the mathematical definition of the covariant derivative uses the Lorenz frame in which the vierbein is diagonalized in the neighborhood. On the other hand, as is shown explicitly above, the vierbein of the local inertial frame is diagonalized only at the point. (Note that the off-diagonal elements vanish at the point since $\sinh a(\tau-\tau_{A})=0$ at $\tau=\tau_{A}$.) Since the covariant derivatives are defined for the Lorentz frame while the Unruh effect is defined for the inertial frame, it seems impossible to examine the Stokes phenomena of the Unruh effect directly (and locally) in terms of the field equations[16]. This mismatch has prevented the local analysis of the Unruh effect for a long time.
@@
Now we consider the physics that observers see in the Unruh effect. As far as the acceleration seen by the observer is constant, the physics seen by the observer is indistinguishable at any time due to the equivalence classes defined for the manifold. (More particularly, the observer feels the same vierbein all the time.) This is purely a mathematical consequence. If a dynamical effect (particle production) is manifested in such a situation, it must be explained by the vierbein of the local inertial frame. Extrapolating the coordinates to infinity and considering a global map as the Bogoliubov transformation is conceptually unacceptable, even if it could be commonly used as a method[6]. In fact, it is known that such methods can lead to unnatural entanglements appearing between regions that should be uncorrelated. To argue the legitimacy of our local computation on manifolds, we must address this issue in this paper.
@@
First of all, consider what happens if the scalar field equations were written down for an accelerating observer. Rindler coordinates are a specific set of coordinates used in the context of special relativity to describe the motion of an observer undergoing constant proper acceleration in flat spacetime. Thus, the Rindler coordinates form a coordinate chart (not a frame) that covers a specific region of Minkowski spacetime known as the Rindler wedge. Following the concept of manifolds, the simplest local inertial system is defined as a tangent space in the neighborhood of the point with zero velocity in Rindler’s coordinates. If one starts with the conventional Rindler metric[6, 12, 41]; [^16] Note that we do not use the metric in our calculations. The reason for showing a metric here is just to explain that the Stokes phenomenon cannot be calculated by the metric.